# Partial Identification Using Neural Causal Models with Continuous Treatment

### v2

@@ 4 Consistency of Neural Causal Partial Identification
Note that our theorem shows that the lower limit of the solution is large than the point $\underline{{F}}^{\widehat{L}_{f}}$, where $\widehat{L}_{f}$ is slightly larger than the original constraint $L_{f}$ we impose on the structural functions. Hence, this point can potentially be slightly smaller than the sharp bound $\underline{{F}}^{L_{f}}$. This worsening is due to the fact that we need to use NNs that satisfy a slightly worse Lipschitz property, to ensure that we have sufficient approximation power. Although Theorem 4.2 does not guarantee $\{F_{n}\}$ converges to a point, it states that $\{F_{n}\}$ may oscillate in the interval $[\underline{F}^{\widehat{L}_{f}},F_{*}]$, which will still give a useful lower bound to ground-truth value $F_{*}$. In particular, if the graph is identifiable, we have $F^{*}=\underline{F}^{\widehat{L}_{f}}$ and $\{F_{n}\}$ converges to $F^{*}$. Also, note that in Theorem 4.2, we use wide NNs rather than deep NN for $\mathcal{F}_{0,n}$ because results in [53] show that wide NNs can approximate Lipschitz functions while controlling their Lipschitz constants (a property that is not yet established for deep NNs). Similar results can be obtained if wide neural nets are used for all components, invoking Theorem 3.5.
As a special case, we leverage Theorem 4.2 for a non-asymptotic rate for the ATE without confounding.
@@
Given two causal models ,^$\mathcal{M},\widehat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$ satisfying 2.3 and 3.2, let their observational distributions be $\nu,\mu$. Suppose the norms of all variables are bounded by $K>0$. If (1) (Overlap) $\nu$ is absolutely continuous with respect to one probability measure $P$ and the density $p_{\nu}\left(t|\text{Pa}(T)=x\right)\geqslant\delta>0$ for x almost surely and $t\in[t_{1},t_{2}]$ and (2) (No Confounding) there is no confounding in the causal graph $\mathcal{G}$, we have
@@
4.1 Experiments
In this section, we examine the performance of our algorithm in two settings. We compare our algorithm with the Autobounds algorithm [14] in a binary IV example in [14] and in a continuous IV model  [^4] The code can be found in https://github.com/Jiyuan-Tan/NeuralPartialID. Since Autobounds can only deal with discrete models, we discretize the continuous model for comparison purposes. The implementation details are provided in the Appendix D. The setting of the first experiment is taken from [14, Section D.1]. This is a binary IV problem and we can calculate the optimal bound using LP [4]. We find that our bound is close to the optimal bound. The second experiment is a general IV example where the treatment is binary but the rest of the variables are continuous. The program that Autobounds solves after discretization contains $\approx 2^{14}$ variables. Even with such a fine discretization, the bound obtained by Autobounds is not tighter than our NCM approach. The details of the structural equations and analysis can be found in Appendix D. We also provide an extra experiment on the counterexample of Proposition C.1 in the appendix to show the effect of Lipschitz regularization.
Setting Algorithm Average Bound Optimal Bound True Value
Binary IV NCM (Ours) [-0.49,0.05] [-0.45, -0.04] -0.31
Autobounds [-0.45,-0.05] [-0.45, -0.04]
General IV NCM (Ours) [2.49,3.24] 3
Autobounds [1.40, 3.48]
Table 1: Experiment results of 2 IV settings. The sample sizes are taken to be $5000$ in each experiment. STD is the standard derivation. The experiments are repeated 10 times for binary IV and 50 times for continuous IV. In all experiments, the bounds given by both algorithms all cover the true values.
Conclusion
@@ B.1 Illustration of how to construct canonical representations and neural architectures
B.3 Proof of results in Section 3.1
Proof of Theorem 3.5.
@@
where $U_{i}$ are i.i.d. uniform random variables on $[0,1]$. By [53, Theorem 2], there exist a deep ReLu network $\widehat{g}_{2}^{i,j}$ of width $\Theta\left(d_{i}^{C}\right)$ and depth $L_{2}$ such that
@@
B.4 Proof of results in Section 3.2
@@
Before we prove Proposition 3.7, we need to introduce some important notions. Let $\{\mathcal{C}_{n}\}$ be a sequence of partitions of the unit cube $[0,1]^{d}$. We say $\{\mathcal{C}_{n}\}$ has property (*) if the following properties are satisfied.
@@
The following lemma is important in the construction of Hilbert curve.
@@
Given a sequence of partitions of the unit cube $\{\mathcal{C}_{n}\}$ that has property (*), there exist a ordering for each partition $\mathcal{C}_{n}$, i.e., $O_{n}:\left\{1,\cdots,2^{nd}\right\}\rightarrow\left\{\overline{i_{1}i_{2}\cdots i_{n}}:i_{j}=0,\cdots,2^{d}-1\right\}$, such that the following two conclusions hold.
@@
Proof of Proposition 3.7.
Without loss of generality, we can assume that $\mathbb{P}$ is supported on $[0,1]^{d}$, vanishes at the boundary and $\mathbb{P}(B)\geqslant C_{1}\lambda(B)$ for some constant $C_{1}>0$ and all measurable sets $B\in[0,1]^{d}$. Suppose we have proven the conclusion for this case. By 3.6, there exists $f\in\mathcal{H}([0,1]^{d},K)\cap\mathcal{F}_{L}([0,1]^{d},K)$ such that $f^{-1}_{\#}\mathbb{P}$ satisfies these conditions. There exists continuous function $\gamma$ such that $f^{-1}_{\#}\mathbb{P}=\gamma_{\#}\lambda$, which implies $\mathbb{P}=(f\circ\gamma)_{\#}\lambda$. In particular, if $\gamma$ is $1/d$-Holder continuous, $f\circ\gamma$ is also $1/d$-Holder continuous because $f$ is Lipschitz continuous.
@@
    Let $A=\prod_{k=1}^{\infty}\left(1+2^{-k}\right)$. We have for any $C\in\mathcal{C}_{n}$, by 3.6
@@
Proof of Theorem 3.8.
The proof is similar to the proof of Theorem 3.5. The difference is that in the first part, we use a deep ReLU network to approximate the Holder continuous curve $\gamma$ instead of wide ReLU network.
Let $\mu_{i}=\mathbb{P}_{C_{i}}/\mathbb{P}(C_{i})$. By 3.4 and 3.6, for each connected component $C_{i}$ of the support, there exists Lipschitz maps $g^{i}_{2}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$ such that
@@
for any measurable set $B\subset[0,1]^{d_{i}^{C}}$. By Proposition 3.7, there exist a $1/d_{i}^{C}$-Holder continuous curve $\gamma_{i}:[0,1]\rightarrow[0,1]^{d_{i}^{C}}$ such that $(\gamma_{i})_{\#}\lambda=(g^{i}_{2})^{-1}_{\#}\mu_{i}$.
By [53, Theorem 2] (take $\omega(x)=C\|x\|^{1/d}$ where $C$ is the Holder continuity constant factor), there exists deep ReLu network $\widehat{g}_{1}^{i,j}$ of width $12$ and depth $L_{1}$ such that
@@
We truncate the neural network at 00 for categorical variables because the propensity functions are required to be non-negative. This truncate operation will not influence the approximation error since $f_{i}$ are non-negative. According to 3.2, $f_{i}$ are Lipschitz continuous. By [53, Theorem 2], we have, if $V_{i}$ is continuous, $\|\widehat{f}_{i}-f_{i}\|_{\infty}\leqslant O(L_{0}^{-1/d_{i}^{\text{in}}})\leqslant O(L_{0}^{-1/d_{\max}^{\text{in}}})$.
@@
$W(P(U_{1},\cdots,U_{n_{U}}),P(\widehat{U}_{1},\cdots,\widehat{U}_{n_{U}}))\leqslant\sum^{n_{U}}_{j=1}W(P(U_{j}),P(\widehat{g}_{j}(Z_{j})))\leqslant O\left(L_{1}^{-2/d_{j}^{U}}+L_{2}^{-2/d_{j}^{U}}+(\tau-\tau\log\tau)\right).$
@@ Appendix C Proof of Consistency
There exists a constant $c>0$ and an SCM $\mathcal{M}^{*}$ satisfying Assumptions 2.3-3.6 with a backdoor causal graph such that there is no unobserved confounding in $\mathcal{M}^{*}$. For any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ with the same causal graph such that $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\epsilon}}(\bm{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
@@
It is easy to see that $\mathcal{M}^{*}$ satisfies Assumption 2.3-3.6. Some calculation gives $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\delta}}(\bm{V}))\leqslant\delta/2$. It is easy to calculate the ATE of the two models.
@@
$\mathcal{P}_{n}=\left\{{P}({\bm{U}}):\bm{U}\ \text{is the latent distribution of}\ \widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n},\mathcal{F}_{2,n},\tau_{n})\right\}.$
In other words, $\mathcal{P}_{n}$ contains all the push-forward measures of the uniform distribution by neural networks. We denote $\Theta_{n}=\widehat{\mathcal{F}}_{V_{1},n}\times\cdots\times\widehat{\mathcal{F}}_{V_{n_{V}},n}\times\times\mathcal{P}_{n}$, where
@@
    [53, Proposition 1] shows that given a $L$-Lipschitz continuous function $f:[0,1]^{m}\rightarrow\mathbb{R}$, there exist $\widehat{f}\in\mathcal{NN}_{m,1}(W,\Theta(\log(m))$ such that $\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/d}\right)$ and $\widehat{f}$ is $\sqrt{m}L$-Lipschitz continuous.  [^6] The original proof does not specify the depth of the neural network. The authors show that the network architectures can be chosen as consisting of $O(W)$ parallel blocks each having the same architecture. Each block is used to realize the function $\phi(x)$ $=(\min(\min_{k\neq s}(x_{k}-x_{s}+1),\min_{k}(1+x_{k}),\min_{k}(1-x_{k})))_{+},\quad x\in\mathbb{R}^{m}.$ This function can be realized by feed-forward network with width $O\left(m^{2}\right)$ and depth $O(\log m)$. For a multivariate function with output dimension $m^{\prime}$ , we can approximate each coordinate individually and get a neural network approximation $\widehat{f}$ that is $\sqrt{mm^{\prime}}L$-Lipschitz continuous and $\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/m}\right)$.

### v3

@@ 4 Consistency of Neural Causal Partial Identification
Note that our theorem shows that the lower limit of the solution is large than the point $\underline{{F}}^{\widehat{L}_{f}}$, where $\widehat{L}_{f}$ is slightly larger than the original constraint $L_{f}$ we impose on the structural functions. Hence, this point can potentially be slightly smaller than the sharp bound $\underline{{F}}^{L_{f}}$. This worsening is due to the fact that we need to use NNs that satisfy a slightly worse Lipschitz property, to ensure that we have sufficient approximation power. Although Theorem 4.2 does not guarantee $\{F_{n}\}$ converges to a point, it states that $\{F_{n}\}$ may oscillate in the interval $[\underline{F}^{\widehat{L}_{f}},F_{*}]$, which will still give a useful lower bound to ground-truth value $F_{*}$. In particular, if the graph is identifiable, we have $F^{*}=\underline{F}^{\widehat{L}_{f}}$ and $\{F_{n}\}$ converges to $F^{*}$. Also, note that in Theorem 4.2, we use wide NNs rather than deep NN for $\mathcal{F}_{0,n}$ because results in Yarotsky2018OptimalAO show that wide NNs can approximate Lipschitz functions while controlling their Lipschitz constants (a property that is not yet established for deep NNs). Similar results can be obtained if wide neural nets are used for all components, invoking Theorem 3.6.
As a special case, we leverage Theorem 4.2 for a non-asymptotic rate for the ATE without confounding. The following proposition proves the Hölder continuity of ATE when there is no confounding.
@@
Given two causal models ,^$\mathcal{M},\widehat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$ satisfying 2.3 and 3.2, let their observational distributions be $\nu,\mu$. Suppose the norms of all variables are bounded by $K>0$. If (1) (Overlap) $\nu$ is absolutely continuous with respect to one probability measure $P$ and the density $p_{\nu}\left(t|\text{Pa}(T)=x\right)\geqslant\delta>0$ for x almost surely and $t\in[t_{1},t_{2}]$ and (2) (No Confounding) there is no hidden confounder in the causal graph $\mathcal{G}$, we have
@@
As a corollary of Proposition 4.3 and Theorem 4.2, we can give a convergence rate for the class of estimands in Theorem 4.2.
@@
5 Experiments
In this section, we examine the performance of our algorithm in two settings. We compare our algorithm with the Autobounds algorithm duarteAutomatedApproachCausal2021 in a binary IV example in duarteAutomatedApproachCausal2021 and in a continuous IV model to calculate bounds for ATE  [^4] The code can be found in https://github.com/Jiyuan-Tan/NeuralPartialID. Since Autobounds can only deal with discrete models, we discretize the continuous model for comparison purposes. The implementation details are provided in the Appendix D. We also provide an extra experiment on the counterexample of Proposition C.1 in the appendix to show the effect of Lipschitz regularization.
5.1 Discrete IV
The setting of the first experiment is taken from (duarteAutomatedApproachCausal2021, , Section D.1). This is a binary IV problem and we can calculate the optimal bound using LP balkeCounterfactualProbabilitiesComputational1994 . The causal graph is shown in Figure 3. We find that our bound is close to the optimal bound. Both algorithms cover the true value in all runs. We could see from Table 1 that the bound given by NCM is slightly worse than Autobounds, but is still close to the optimal bound. Besides, the avgerage Autobounds bound does not cover the optimal bound in this example.
ZTYU
Algorithm Average Bound SD of length Optimal Bound True Value
NCM (Ours) [-0.49,0.05] 0.05 [-0.45, -0.04] -0.31
Autobounds (duarteAutomatedApproachCausal2021 ) [-0.45,-0.05] 0.02 [-0.45, -0.04] -0.31
Table 1: The sample size is taken to be $5000$. We average over 10 runs with different random seeds. SD is short for the standard deviation. In all runs, the bounds given by the two algorithms all cover the true value.
5.2 Continuous IV
$E_{Y}=E^{\lambda}_{1},$ $\ U=E^{\lambda}_{2},Z=E^{\lambda}_{3},$
$P(T=1|Z)$ $=(1.5+Z+0.5U)/3,$
$Y$ $=3T-1.5TU+U+E_{Y},$
Table 2 demonstrates the results. While both algorithms cover the true ATE well, we can see that NCM gives much tighter bounds on average. The main reason may be that the discretized problem does not approximate the original problem well enough. It is possible that a larger discretized parameter $k$ can give a better bound, but since the size of the polynomial problem grows exponentially with $k$, the optimization problem may be intractable for large $k$. On the contrary, NCM does not suffer from computational difficulties.
Algorithm Average Bound SD of length Success Rate True Value
NCM (Ours) [2.49,3.24] 0.49 50/50 3
Autobounds (duarteAutomatedApproachCausal2021 ) [1.40, 3.48] 0.26 50/50 3
Table 2: We take a sample size of 5000. We randomly choose 10 $\lambda\sim\text{Unif}(0,1)$ and get 10 models $\mathcal{M}^{\lambda_{i}}$. For each model $\mathcal{M}^{\lambda_{i}}$, we run the two algorithms 5 times. The success rate is the number of times when the obtained bounds cover the true ATE divided by the total number of experiments. SD is short for the standard deviation.
6 Conclusion
@@ B.1 Illustration of how to construct canonical representations and neural architectures
B.3 Proof of results in Section 3.1: Approximating with Wide NN
Proof of Theorem 3.6 Part I.
@@
where $U_{i}$ are i.i.d. uniform random variables on $[0,1]$. By [52, Theorem 2], there exist a deep ReLu network $\widehat{g}_{2}^{i,j}$ of width $\Theta\left(d_{i}^{C}\right)$ and depth $L_{2}$ such that
@@
B.4 Proof of results in Section 3.1: Approximating with Deep NN
@@
We first start with an important proposition.
Proposition B.2.
Let $\lambda$ be the Lebesgue measure on $[0,1]$. Given any probability measure $\mathbb{P}$ that satisfies 3.5, there exists a continuous curve $\gamma:[0,1]\rightarrow\text{supp}(\mathbb{P})$ such that $\gamma_{\#}\lambda=\mathbb{P}$. Furthermore, if $d\geqslant 1$, $\gamma$ is $1/d$-Hölder continuous.
Before we prove Proposition B.2, we need to introduce some important notions. Let $\{\mathcal{C}_{n}\}$ be a sequence of partitions of the unit cube $[0,1]^{d}$. We say $\{\mathcal{C}_{n}\}$ has property (*) if the following properties are satisfied.
@@
The following lemma is important in the construction of the Hilbert curve.
@@
Given a sequence of partitions of the unit cube $\{\mathcal{C}_{n}\}$ that has property (*), there exist a ordering for each partition $\mathcal{C}_{n}$, i.e.,
$O_{n}:\left\{1,\cdots,2^{nd}\right\}\rightarrow\left\{\overline{i_{1}i_{2}\cdots i_{n}}:i_{j}=0,\cdots,2^{d}-1\right\},$
such that the following two conclusions hold.
@@
Proof of Proposition B.2.
Without loss of generality, we can assume that $\mathbb{P}$ is supported on $[0,1]^{d}$, vanishes at the boundary and $\mathbb{P}(B)\geqslant C_{1}\lambda(B)$ for some constant $C_{1}>0$ and all measurable sets $B\in[0,1]^{d}$. Suppose we have proven the conclusion for this case. By 3.5, there exists $f\in\mathcal{H}([0,1]^{d},K)\cap\mathcal{F}_{L}([0,1]^{d},K)$ such that $f^{-1}_{\#}\mathbb{P}$ satisfies these conditions. There exists continuous function $\gamma$ such that $f^{-1}_{\#}\mathbb{P}=\gamma_{\#}\lambda$, which implies $\mathbb{P}=(f\circ\gamma)_{\#}\lambda$. In particular, if $\gamma$ is $1/d$-Holder continuous, $f\circ\gamma$ is also $1/d$-Holder continuous because $f$ is Lipschitz continuous.
@@
    Let $A=\prod_{k=1}^{\infty}\left(1+2^{-k}\right)$. We have for any $C\in\mathcal{C}_{n}$, by 3.5
@@
Proof of Theorem 3.6 Part II.
The proof is similar to the proof of Theorem 3.6 Part I. The difference is that in the first part, we use a deep ReLU network to approximate the Holder continuous curve $\gamma$ instead of wide ReLU network.
Let $\mu_{i}=\mathbb{P}_{C_{i}}/\mathbb{P}(C_{i})$. By 3.4 and 3.5, for each connected component $C_{i}$ of the support, there exists Lipschitz maps $g^{i}_{2}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$ such that
@@
for any measurable set $B\subset[0,1]^{d_{i}^{C}}$. By Proposition B.2, there exist a $1/d_{i}^{C}$-Holder continuous curve $\gamma_{i}:[0,1]\rightarrow[0,1]^{d_{i}^{C}}$ such that $(\gamma_{i})_{\#}\lambda=(g^{i}_{2})^{-1}_{\#}\mu_{i}$.
By [52, Theorem 2] (take $\omega(x)=C\|x\|^{1/d}$ where $C$ is the Holder continuity constant factor), there exists deep ReLu network $\widehat{g}_{1}^{i,j}$ of width $12$ and depth $L_{1}$ such that
@@
We truncate the neural network at 00 for categorical variables because the propensity functions are required to be non-negative. This truncate operation will not influence the approximation error since $f_{i}$ are non-negative. According to 3.2, $f_{i}$ are Lipschitz continuous. By [52, Theorem 2], we have, if $V_{i}$ is continuous, $\|\widehat{f}_{i}-f_{i}\|_{\infty}\leqslant O(L_{0}^{-1/d_{i}^{\text{in}}})\leqslant O(L_{0}^{-1/d_{\max}^{\text{in}}})$.
@@
$W(P(U_{1},\cdots,U_{n_{U}}),P(\widehat{U}_{1},\cdots,\widehat{U}_{n_{U}}))$ $\leqslant\sum^{n_{U}}_{j=1}W(P(U_{j}),P(\widehat{g}_{j}(Z_{j})))$
$\leqslant O\left(\sum^{n_{U}}_{j=1}(L_{1}^{-2/d_{j}^{U}}+L_{2}^{-2/d_{j}^{U}})+(\tau-\tau\log\tau)\right).$
@@ Appendix C Proof of Consistency
There exists a constant $c>0$ and an SCM $\mathcal{M}^{*}$ satisfying Assumptions 2.3-3.5 with a backdoor causal graph such that there is no unobserved confounding in $\mathcal{M}^{*}$. For any $\epsilon>0$, there exists an SCM $\mathcal{M}_{\epsilon}$ with the same causal graph such that $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\epsilon}}(\bm{V}))\leq\epsilon$ and $|\text{ATE}_{\mathcal{M}^{*}}-\text{ATE}_{\mathcal{M}_{\epsilon}}|>c$.
@@
It is easy to see that $\mathcal{M}^{*}$ satisfies Assumption 2.3-3.5. Some calculation gives $W(P^{\mathcal{M}^{*}}(\bm{V}),P^{\mathcal{M}_{\delta}}(\bm{V}))\leqslant\delta/2$. It is easy to calculate the ATE of the two models.
@@
$\mathcal{P}_{n}=\left\{{P}({\bm{U}}):\bm{U}\ \text{is the latent distribution of}\ \widehat{\mathcal{M}}\in\text{NCM}_{\mathcal{G}}(\mathcal{F}_{0,n},\mathcal{F}_{1,n})\right\}.$
In other words, $\mathcal{P}_{n}$ contains all the push-forward measures of the uniform distribution by neural networks. We denote $\Theta_{n}=\widehat{\mathcal{F}}_{V_{1},n}\times\cdots\times\widehat{\mathcal{F}}_{V_{n_{V}},n}\times\mathcal{P}_{n}$, where
@@
    [52, Proposition 1] shows that given a $L$-Lipschitz continuous function $f:[0,1]^{m}\rightarrow\mathbb{R}$, there exist $\widehat{f}\in\mathcal{NN}_{m,1}(W,\Theta(\log(m))$ such that $\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/d}\right)$ and $\widehat{f}$ is $\sqrt{m}L$-Lipschitz continuous.  [^6] The original proof does not specify the depth of the neural network. The authors show that the network architectures can be chosen as consisting of $O(W)$ parallel blocks each having the same architecture. Each block is used to realize the function $\phi(x)$ $=(\min(\min_{k\neq s}(x_{k}-x_{s}+1),\min_{k}(1+x_{k}),\min_{k}(1-x_{k})))_{+},\quad x\in\mathbb{R}^{m}.$ This function can be realized by feed-forward network with width $O\left(m^{2}\right)$ and depth $O(\log m)$. For a multivariate function with output dimension $m^{\prime}$ , we can approximate each coordinate individually and get a neural network approximation $\widehat{f}$ that is $\sqrt{mm^{\prime}}L$-Lipschitz continuous and $\|f-\widehat{f}\|_{\infty}\leqslant O\left(W^{-1/m}\right)$.